2-137. The following circuit operates if and only if there is a path of functional devices from left to right. The probability each device functions is as shown. Assume that the probabil- ity that a device functions does not depend on whether or not other devices are functional. What is the probability that the circuit operates? 0.9 _ 0.9- 0.8- 0.95 0.95- 0.9
Answer #1

here there are 3 parallel blocks

probbability that one parallel bock works =1-P(both of the component of parallel block does not work)

=1-(1-0.9)*(1-0.9)=0.99

therefore probability that circuit operates =0.99*0.99*0.99=0.9703